Jarmon:

Suppose that y varies inversely with x. Use the information to find k, and then choose the equation of variation. X= 28 when y=2

Nnesha:

Same concept: `y varies inversely with x' is y =k/x replace `y` with 2 and x with 28 \[2=\frac{k}{28}\] can you solve for y ??

Nnesha:

for k** ugh

Jarmon:

Nah

Nnesha:

why not ?? multiply both sides by 28 what do you get ?

Jarmon:

56

Nnesha:

yes now replace k with 56 y=56/x that's it